A California firm is planning to use a lockbox system to speed up collections from its customers located on the East Coast. A Philadelphia-area bank will provide this service for an annual fee of $6500 plus 10 cents per transaction. The estimated reduction in collection and processing time is one day. If the average customer payment in this region is $2900, how many customers each day, on average, are needed to make the system profitable for the firm? The firm can invest in Treasury bills, currently yielding 4% per year, and there are 365 days per year.
Answer #1

Total fee = $6500 + 0.1 * N *365

Per day yield of treasury bills = (1+ 4%)1/365 -1 = 0.00010746

Payments received for the customers = 2900* N

The system is profitable when NPV is $0.

0 = 2900 *N- 0.1*N / 0.00010746 -6500/0.04

N = 82.51 customers

Any doubt please comment.